OpenStudy (anonymous):

PLEASE HELP! Solve the equation 8^9x=16

OpenStudy (anonymous):

????

OpenStudy (mathmate):

Do you mean \( 8^{9x}=16\) ? \((2^3)^{9x} = 2^4\) \(2^{27x}=2^4\) 27x = 4 Can you continue to find x?

OpenStudy (anonymous):

It just states to solve the given equation and I am working on logarithms. My answer choices are a. 2.7726 c. 2.0794 b. 0.1481 d. 1.9685

OpenStudy (mathmate):

If you continue my calculations, the answer for x is one of your choices!

OpenStudy (anonymous):

Which one? Please help i couldn't get it

OpenStudy (mathmate):

Can you find x from the following equation: 27x = 4

OpenStudy (anonymous):

Got it! Thanks a lot

OpenStudy (mathmate):

You're welcome! :)